The speed of a stream is 4mph. A boat travels 8 miles upstream in the same time it takes to travel 16 miles downstream. What is the speed of the boat in the water?
Answer by mananth(16949)

let the speed of boat in still water be x mph
speed of current 4 mph
..
speed upstream = x-4 mph
speed downstream = x+4
..
distance = 8 miles upstream
time upstream = 8/x-4
..
time downstream = 16/x+4)
..
both the times are same
..
8/(x-4) = 16/(x+4)
8(x+4)=16(x-4)
8x+32=16x-64
16x-8x=64+32
8x=96
x= 12 mph speed of boat in still water